OpenStudy (anonymous):

The mean for average temperature in Buffalo ,New york , is 47.5°. The temperture fluctuates 23.5° above and below the mean temperture . If t=1 present January ,the phase shift of the sine function is 4 a) Write the funtion for the average monthly tempperature in Buffalo b) According to your modle ,what is the average tenperature in March ? c) According to your modle ,what is the average tenperature in August ?

OpenStudy (anonymous):

\[a \sin( bx +c ) + d \] a influences amplitude ( a = 1 => -1 < sin < 1 ) b influences frequency ( b = 1 => i cycle takes 2 pi ) c phase shift d vertical translation So what you need is the following mean is 47.5 so d = 47.5 23.5 fluctiation means a = 23.5 The last bit is a bit confusing, but i'll give it a try: phase shift is 4 which makes c 4 now we want it to have a frequency of 12 months (at least i think that's what's needed here) we need b to be pi/6 Conclusion: \[T(x) = 23.5 \sin( { \pi \over 6}x +4) +47.5 (x = months)\] for quetions b and c simply insert the monthnumber in this function

OpenStudy (anonymous):

What's tricky is the monthnumber, 1 represents the 31st of jan in this formula, not the entire month of jan. since it's a continuous function.